### v1

@@ 6 Discussion and Future Research Directions
    Integrating recommendation systems with developer workflow tools involves embedding recommendation functionalities directly into software development tools and environments commonly used by developers, such as IDEs, version control systems, project management tools, or communication platforms. For example, Study S1 suggests integrating the proposed recommendation system with Github actions or bots. Study S3 developed an open source plugin for some of JetBrains’ open source IDEs such as IntelliJ IDEA, CLion, etc.
    Integrating the recommendation system with developer workflow tools streamlines the task discovery and assignment process, enabling developers to access relevant tasks directly within their existing development environment without needing to switch between different tools or platforms.
    However, different development tools and environments have varied architectures and APIs that make the integration process complex [63]. For example, integrating with Eclipse IDE might differ significantly from integrating with Visual Studio IDE [64]. This can be addressed by utilizing a plugin-based architecture where the recommendation functionalities are provided as plugins or extensions for different IDEs and tools [64] [65]. Further, maintaining compatibility with continuously evolving tools and the latest versions of the IDEs can be challenging. Another key challenge of integrating recommendation systems with developer workflow tools is that recommendation systems can consume significant computational resources, potentially slowing down the development environment [66].
    We suggest researchers integrate their recommendation systems with developer workflow tools.
@@
    Interactive recommendation systems enable users to actively engage with the recommendation process by providing feedback, preferences, and additional input [67]. Tailoring recommendations to individual preferences increases the relevancy of the recommendations. Continuous feedback may help refine and improve the recommendation algorithms over time. Hence, such interaction enhances user satisfaction and improves the user experience by adapting to changing preferences over time [68]. For example, Study S3 currently allows users (specifically Eclipse developers) to provide feedback through likes, dislikes, or snoozes (reminders) for issues. They utilize this feedback to hide certain issues in case of dislikes and to re-rank the recommendation list based on likes. In the future, they intend to leverage this data to automatically enhance and adjust the recommendation model.
@@
    Additionally, the development of Large Language Models (LLMs), like ChatGPT and GPT-4, has transformed the fields of Natural Language Processing (NLP) and Artificial Intelligence (AI). These models exhibit exceptional proficiency in fundamental tasks of language understanding and generation, alongside notable generalization abilities and reasoning skills [72]. Hence, LLMs can be employed to provide interactive recommendations as those excel at processing and understanding natural language input and leverage context from user queries and previous interactions to generate more relevant and personalized recommendations [73].
@@
Integrating recommendation systems with developer workflow tools
@@ 7 Conclusion
This SLR has provided a detailed analysis of software task recommendations to software practitioners in crowdsourcing platforms. Our analysis equips researchers with a comprehensive understanding of existing research in the domain, including the existing methods, benefits, and limitations of those methods, different task types recommended, use of human factors in recommendations, popular crowdsourcing platforms, and the features of those platforms used to extract data. Further, we highlighted the key areas for future research. For the SLR, We adhered to the systematic process outlined by Kitchenham et al [13]. We selected 63 highly relevant primary studies for the SLR after a rigorous filtration process. Our main findings indicate that the number of studies on crowdsourced software engineering task recommendations significantly increased over the past couple of years, and we anticipate that this trend will continue. The identified future directions for further research in this domain include: (1) More studies are needed to explore the effect of different human factors such as the personality of software practitioners, motivation, and career goals on CSE; (2) limited attention has been paid to transferring knowledge from other related platforms, which can be beneficial in addressing the “cold start” problem; (3) Investigation is needed to introduce generalizable platform-independent recommendation systems; (4) Need to conduct comprehensive further evaluation on existing systems and need establish standardized evaluation metrics and bench-marking; (5) Investigation needed to integrate the recommendation system with developer workflow tools; (6) Can develop existing systems to interactive recommendation systems with the help of large language models which enable users to actively engage with the recommendation process.

### v2

@@ 6 Discussion and Future Research Directions
    Integrating recommendation systems with development workflow tools involves embedding recommendation functionalities directly into software development tools and environments commonly used by developers, such as IDEs, version control systems, project management tools, or communication platforms. For example, Study S1 suggests integrating the proposed recommendation system with Github actions or bots. Study S3 developed an open source plugin for some of JetBrains’ open source IDEs such as IntelliJ IDEA, CLion, etc.
    Integrating the recommendation system with development workflow tools streamlines the task discovery and assignment process, enabling developers to access relevant tasks directly within their existing development environment without needing to switch between different tools or platforms. Further, current adaptive IDE systems, such as the one proposed by Schmidmaier et al. [73], can track the task context, developer’s expertise level, and interaction patterns directly from the IDE. In crowdsourced development, contributors are diverse and work remotely. Unlike traditional teams, where roles and expertise are known, crowdsourcing platforms lack direct visibility into a developer’s real-time interests and skill level. By leveraging real-time information from IDEs, recommendation systems can refine their recommendations to better align with the developer’s context. For instance, if the IDE detects that a developer is working on a machine learning project and has an advanced expertise level, the system can prioritize recommending complex ML-related crowdsourced tasks.
    However, different development tools and environments have varied architectures and APIs that make the integration process complex [74]. For example, integrating with Eclipse IDE might differ significantly from integrating with Visual Studio IDE [75]. This can be addressed by utilizing a plugin-based architecture where the recommendation functionalities are provided as plugins or extensions for different IDEs and tools [75] [76]. Further, maintaining compatibility with continuously evolving tools and the latest versions of the IDEs can be challenging. Another key challenge of integrating recommendation systems with development workflow tools is that recommendation systems can consume significant computational resources, potentially slowing down the development environment [77].
    We suggest researchers integrate their recommendation systems with development workflow tools.
@@
    Interactive recommendation systems enable users to actively engage with the recommendation process by providing feedback, preferences, and additional input [78]. In the context of CSE, where contributors vary in expertise, availability, and motivations, interactive recommendations can help match developers with tasks more effectively. Tailoring recommendations to individual preferences increases the relevancy of the recommendations. Continuous feedback may help refine and improve the recommendation algorithms over time. Hence, such interaction enhances user satisfaction and improves the user experience by adapting to changing preferences over time [79, 80]. For example, Study S3 currently allows users (specifically Eclipse developers) to provide feedback through likes, dislikes, or snoozes (reminders) for issues. They utilize this feedback to hide certain issues in case of dislikes and to re-rank the recommendation list based on likes. In the future, they intend to leverage this data to automatically enhance and adjust the recommendation model.
@@
    Additionally, the development of Large Language Models (LLMs), like ChatGPT and GPT-4, has transformed the fields of Natural Language Processing (NLP) and Artificial Intelligence (AI). These models exhibit exceptional proficiency in fundamental tasks of language understanding and generation, alongside notable generalization abilities and reasoning skills [84]. Study S64 mentioned in the future works, integrating LLMs like GPT with their recommendation system can significantly improve recommendation accuracy by capturing complex linguistic patterns and semantics. LLMs could assist in dynamically refining task recommendations by processing natural language queries, extracting intent from developer interactions, and suggesting relevant tasks that align with evolving skills and interests. Hence, LLMs can be employed to provide interactive recommendations as those excel at processing and understanding natural language input and leverage context from user queries and previous interactions to generate more relevant and personalized recommendations [85]. LLMs can help address the cold-start problem by leveraging limited data to generate personalized, context-aware recommendations. Their ability to process customizable prompts in chat-based systems enhances user engagement. For example, the LLM-based recommendation system in [86] allows users to express their preferences and intents in natural language. This approach outperforms traditional methods relying solely on user-item interactions. Given their effectiveness in tackling data sparsity and improving efficiency, language models have become a promising approach in advancing recommendation systems in both academia and industry [79].
@@
Integrating recommendation systems with development workflow tools
@@ 7 Conclusion
This SLR has provided a detailed analysis of software task recommendations to software practitioners in crowdsourcing platforms. Our analysis equips researchers with a comprehensive understanding of existing research in the domain, including the existing methods, benefits, and limitations of those methods, different task types recommended, use of human factors in recommendations, popular crowdsourcing platforms, and the features of those platforms used to extract data. Further, we highlighted the key areas for future research. For the SLR, We adhered to the systematic process outlined by Kitchenham et al [13]. We selected highly relevant primary studies for the SLR after a rigorous filtration process. Our main findings indicate that the number of studies on crowdsourced software engineering task recommendations significantly increased over the past couple of years, and we anticipate that this trend will continue. The identified future directions for further research in this domain include: (1) More studies are needed to explore the effect of different human factors such as the personality of software practitioners, motivation, and career goals on CSE; (2) limited attention has been paid to transferring knowledge from other related platforms, which can be beneficial in addressing the “cold start” problem; (3) Investigation is needed to introduce generalizable platform-independent recommendation systems; (4) Need to conduct comprehensive further evaluation on existing systems and need establish standardized evaluation metrics and bench-marking; (5) Investigation needed to integrate the recommendation system with development workflow tools; (6) Can develop existing systems to interactive recommendation systems with the help of large language models which enable users to actively engage with the recommendation process.
@@ Appendix B Quality Assesment Scores
[143] S64 CodeCompass: NLP-Driven Navigation to Optimal Repositories International Conference on Pervasive Computing and Social Networking 2024
[144] S65 Multi-objective optimization and integrated indicator-driven two-stage project recommendation in time-dependent software ecosystem Information and Software Technology 2024
Table 16: Quality assessment scores for primary studies
ID QA1 QA2 QA3 QA4 QA5 ID QA1 QA2 QA3 QA4 QA5
S1 5 5 4 5 2 S34 4 3 3 3 3
S2 5 4 4 5 4 S35 4 3 2 4 2
S3 4 5 3 3 3 S36 3 3 2 3 2
S4 3 2 2 3 1 S37 3 3 3 1 1
S5 3 3 4 4 2 S38 3 3 3 3 3
S6 3 2 3 4 2 S39 4 4 3 3 3
S7 5 3 2 3 1 S40 3 2 4 3 2
S8 4 2 4 4 2 S41 3 1 2 1 1
S9 4 3 4 4 4 S42 4 3 3 3 4
S10 3 2 3 3 2 S43 4 2 4 4 3
S11 3 4 3 4 4 S44 3 4 3 3 2
S12 4 3 3 4 4 S45 4 2 3 1 1
S13 3 2 3 3 3 S46 3 3 4 4 2
S14 3 3 4 4 2 S47 4 5 2 3 1
S15 4 4 4 3 3 S48 4 4 3 4 3
S16 3 2 3 3 2 S49 4 3 3 3 1
S17 3 3 3 4 3 S50 3 3 4 4 4
S18 3 3 2 3 1 S51 4 3 3 3 2
S19 2 3 1 3 1 S52 2 4 3 1 1
S20 4 4 3 3 3 S53 4 3 4 4 5
S21 4 3 2 3 2 S54 3 4 3 4 2
S22 3 4 3 4 4 S55 3 4 3 3 2
S23 4 3 3 4 4 S56 4 3 3 3 3
S24 4 3 3 4 3 S57 3 2 3 3 1
S25 3 3 3 3 3 S58 2 2 2 3 1
S26 2 3 3 1 1 S59 4 3 3 4 3
S27 4 3 4 4 2 S60 3 3 3 3 1
S28 3 3 5 3 S61 4 3 4 3 1
S29 3 3 4 4 2 S62 3 3 3 4 1
S30 3 3 2 2 2 S63 3 3 2 2 1
S31 3 3 4 4 4 S64 3 3 3 3 2
S32 3 4 2 3 1 S65 3 3 4 4 1
S33 4 4 3 3 3